OpenStudy (anonymous):

Solve: u2 + 2u - 3 = 0

OpenStudy (callisto):

u^2 + 2u - 3 = 0 (u+3)(u-1) =0 u+3 =0 or u-1=0 Can you solve it from here?

OpenStudy (anonymous):

i dont know how

OpenStudy (callisto):

Well. for u+3 =0 u+3 =0 -3 from both side, what do you get?

OpenStudy (anonymous):

he get that he's dumb !! LOL